# Variational Inference for Sparse Poisson Regression

### v8

@@ Abstract
We have utilized the non-conjugate Variational Bayesian (VB) method for the problem of the sparse Poisson regression model. To provide an approximate conjugacy in the model, the likelihood is approximated by a quadratic function, which provides the conjugacy of the approximation component with the Gaussian prior on the regression coefficient. Three sparsity-enforcing priors are used for this problem. The proposed models are compared with each other and two frequentist sparse Poisson methods (LASSO and SCAD) to evaluate the estimation, prediction, and sparsity performance of the proposed methods. In a simulation study, the proposed VB methods closely approximate the posterior parameter distribution while achieving significantly faster computation than benchmark MCMC methods. Using several benchmark count response data sets, the prediction performance of the proposed methods is evaluated in real-world applications.
@@ 1 Introduction
Although MCMC is the most widely used approach for approximating the posterior distribution in Bayesian inference, it incurs a high computational cost, particularly when the dimensionality of the regression parameters is large. There are a variety of methods for approximation of the posterior distribution with a lower computational cost, including integrated nested Laplace approximation (INLA- Rue et al., 2010), mean-field variational Bayes (VB- Blei et al., 2017), expectation propagation (EP- Minka, 2001, 2013), and message passing (MP- Winn et al., 2005). The mean-field VB is the most frequently used method in this category. The VB method is applied in many problems, especially in Bayesian regression models (see Šmídl and Quinn, 2006; Minka, 2000; Wand et al., 2011; Broderick et al., 2013; Drugowitsch, 2013; Goldsmith et al., 2011; Pham et al., 2013; Titsias and Lázaro-Gredilla, 2011, and the references therein). The VB method is also applied to the sparse linear regression models (Ormerod et al., 2017; Ray and Szabó, 2022; Titsias and Lázaro-Gredilla, 2011; Huang et al., 2016).
@@
The rest of the paper is organized as follows. In Section 2, the mean-field variational Bayes method is introduced. Three variational Bayesian sparse Poisson regression models are then proposed in Section 3. The derivation of the components, as well as the evidence lower bound (ELBO) for these methods, is given in Appendix A. In Section 4, a hard thresholding method is introduced to derive the final sparse estimates of the regression coefficients. Section 5 introduces the posterior predictive density function for predicting the response variable based on new covariate samples. Section 6 presents a simulation study for evaluating the estimation, prediction, and sparsity performance of the three proposed methods and comparing them with two frequentist sparse Poisson regression methods (LASSO and SCAD Poisson regression models). A single replication of the simulation study is also used for the computation of the accuracy of each approximation component compared to its corresponding benchmark MCMC method, and also a visual comparison of the MCMC empirical posteriors with their approximated counterparts. Some of the plots and additional analysis are given in supplementary material for the sake of brevity. Finally, in Section 7, the test set prediction performance of the proposed methods is compared with that of the frequentist alternatives, using 6 benchmark count response data sets. Some concluding remarks are given in Section 8.
@@ 2 Mean field variational Bayes
$q({\boldsymbol{\theta}})=\prod_{j=1}^{M}q({\boldsymbol{\theta}}_{j}).$
@@ 3.1 Laplace prior
Since, the equality holds in (4) when $\xi=x$, we let $\xi_{i}={\rm E}_{q}(\mathbf{X}_{i}{\boldsymbol{\beta}})=\mathbf{X}_{i}{\boldsymbol{\mu}}^{L}_{{\boldsymbol{\beta}}}$, for $i=1,\ldots,n$, where ${\boldsymbol{\mu}}^{L}_{{\boldsymbol{\beta}}}$ is given in (6). Let $\boldsymbol{\xi}=(\xi_{1},\ldots,\xi_{n})^{\top}$, and define $M_{\boldsymbol{\xi}}=e^{\xi}(1-\xi)$ and $S_{X}^{\boldsymbol{\xi}}=\sum_{i=1}^{n}e^{\xi_{i}}\mathbf{X}_{i}\mathbf{X}_{i}^{\top}$.
@@
${\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(L)}=\left[S_{X}^{\boldsymbol{\xi}}+\text{diag}{\rm E}({\boldsymbol{\tau}}^{-1})\right]^{-1},\quad{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(L)}={\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(L)}\mathbf{X}^{\top}(\mathbf{y}-M_{\boldsymbol{\xi}}).$ (6)
Let $D^{L}_{{\boldsymbol{\beta}}}={\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(L)}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(L)}^{\top}+{\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(L)}$. Furthermore,
$q({\eta})=\text{Gamma}({\eta};{p+\nu-1},\delta+\frac{1}{2}\sum_{j=1}^{p-1}{\rm E}(\tau_{j})),$
and ${\rm E}({\eta})=\frac{{p+\nu-1}}{\delta+\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}(\tau_{j})}$, for $j=1,\ldots,p$,
$q(\tau_{j})=\text{Gen-Inv-Gauss}\Big{(}\tau_{j};\frac{1}{2},{\rm E}({\eta}),{\rm E}(\beta_{j}^{2})\Big{)},$
@@
${\rm E}(\tau_{j})=\frac{\sqrt{{\rm E}(\beta_{j}^{2})}K_{3/2}(\sqrt{{\rm E}({\eta}){\rm E}(\beta_{j}^{2})})}{\sqrt{{\rm E}({\eta})}K_{1/2}(\sqrt{{\rm E}({\eta}){\rm E}(\beta_{j}^{2})})},$
${\rm E}(\tau_{j}^{-1})=\frac{\sqrt{{\rm E}({\eta})}K_{3/2}(\sqrt{{\rm E}({\eta}){\rm E}(\beta_{j}^{2})})}{\sqrt{{\rm E}(\beta_{j}^{2})}K_{1/2}(\sqrt{{\rm E}({\eta}){\rm E}(\beta_{j}^{2})})}-\frac{1}{{\rm E}(\beta_{j}^{2})},$
@@
${\rm E}(\log\tau_{j})=\log\sqrt{\frac{{\rm E}(\beta_{j}^{2})}{{\rm E}({\eta})}}+\frac{\partial}{\partial t}\log K_{t}(\sqrt{{\rm E}({\eta}){\rm E}(\beta_{j}^{2})})|_{t=1/2}.$
@@
$q(\tau_{0})=\text{Inv-Gamma}(\tau_{0};1,\frac{1}{2}d_{00}^{L}+{\rm E}(a^{-1})),$
where ${d_{jj}^{L}}={\rm E}(\beta_{j}^{2})=(D^{L}_{{\boldsymbol{\beta}}})_{jj}$, and thus ${\rm E}(\tau_{0}^{-1})=\dfrac{1}{\frac{1}{2}d_{00}+{\rm E}(a^{-1})}$ and ${\rm E}(\log\tau_{0})=\log(\frac{1}{2}d_{00}+{\rm E}(a^{-1}))-\psi(1)$, where $\psi(\cdot)$ is the digamma function, and
$q(a)=\text{Inv-Gamma}(a;1,{\rm E}(\tau_{0}^{-1})+A^{-1}),$
and therefore ${\rm E}(a^{-1})=\dfrac{1}{{\rm E}(\tau_{0}^{-1})+A^{-1}}$ and ${\rm E}(\log a)=-\psi(1)+\log({\rm E}(\tau_{0}^{-1})+A^{-1}))$.
@@
$\log p(\mathbf{y};q)$ $\approx{{\rm E}_{q}(\log\tilde{p}(\mathbf{y}|{\boldsymbol{\beta}},\boldsymbol{\xi}))+{\rm E}_{q}(\log p({\boldsymbol{\beta}}|\tau))+{\rm E}_{q}(\log p({\eta}))\allowdisplaybreaks}$
${\quad+{\rm E}_{q}(\log p(\tau_{0}|a))+{\rm E}_{q}(\log p(a))-{\rm E}_{q}(\log q({\boldsymbol{\beta}}))\allowdisplaybreaks}$
${\quad-{\rm E}_{q}(\log q({\eta}))-{\rm E}_{q}(\log q(\tau_{0}))-{\rm E}_{q}(\log q(\tau))-{\rm E}_{q}(\log q(a))\allowdisplaybreaks}$
$=-M_{\boldsymbol{\xi}}^{T}(1+\mathbf{X}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(L)})-\frac{1}{2}\sum_{i=1}^{n}\xi_{i}^{2}e^{\xi_{i}}-\frac{1}{2}tr(S_{X}^{\boldsymbol{\xi}}D_{{\boldsymbol{\beta}}}^{L})+\mathbf{y}^{\top}\mathbf{X}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(L)}\allowdisplaybreaks$
$\quad-\frac{1}{2}{\rm E}(\log\tau_{0})-\frac{1}{2}{\rm E}(\tau_{0}^{-1}){\rm E}(\beta_{0}^{2})-\dfrac{\delta({p+\nu-1})}{\delta+\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}(\tau_{j})}\allowdisplaybreaks$
$\quad-2\log(A^{-1}+{\rm E}(\tau_{0}^{-1}))-\frac{1}{2}\log(d_{00}^{L}/2+{\rm E}(a^{-1}))\allowdisplaybreaks$
$\quad-{\rm E}(a^{-1}){\rm E}(\tau_{0}^{-1})-\dfrac{A^{-1}}{A^{-1}+E(\tau_{0}^{-1})}+\frac{1}{2}\log|{\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(L)}|\allowdisplaybreaks$
$\quad-\frac{1}{4}{\sum_{j=1}^{p-1}}\frac{{\rm E}({\eta})}{{d_{jj}^{L}}}+{\sum_{j=1}^{p-1}}\log K_{1/2}(\sqrt{{\rm E}({\eta}){d_{jj}^{L}}})+\text{Const.}$ (7)
@@
  $\bullet$ update ${\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(L)}\leftarrow\left[S_{X}^{\boldsymbol{\xi}}+\text{diag}{\rm E}({\boldsymbol{\tau}}^{-1})\right]^{-1},$ and ${\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(L)}\leftarrow{\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(L)}\mathbf{X}^{\top}(\mathbf{y}-M_{\boldsymbol{\xi}}),$
@@
  $\bullet$ update ${\rm E}(\beta_{j}^{2})={d_{jj}^{L}}\leftarrow{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(L)j}^{2}+{\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(L)jj}$, for $j=0,\ldots,p-1$,
  $\bullet$ update ${\rm E}({\eta})\leftarrow\frac{{p+\nu-1}}{\delta+\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}(\tau_{j})}$
  $\bullet$ update ${\rm E}(\tau_{j})\leftarrow\frac{\sqrt{{\rm E}(\beta_{j}^{2})}K_{3/2}(\sqrt{{\rm E}({\eta}){\rm E}(\beta_{j}^{2})})}{\sqrt{{\rm E}({\eta})}K_{1/2}(\sqrt{{\rm E}({\eta}){\rm E}(\beta_{j}^{2})})},$ for $j=1,\ldots,p-1$,
  $\bullet$ update ${\rm E}(\tau_{j}^{-1})\leftarrow\frac{\sqrt{{\rm E}({\eta})}K_{3/2}(\sqrt{{\rm E}({\eta}){\rm E}(\beta_{j}^{2})})}{\sqrt{{\rm E}(\beta_{j}^{2})}K_{1/2}(\sqrt{{\rm E}({\eta}){\rm E}(\beta_{j}^{2})})}-\frac{1}{{\rm E}(\beta_{j}^{2})},$ for $j=1,\ldots,p-1$,
  $\bullet$ update ${\rm E}(\tau_{0}^{-1})\leftarrow\dfrac{1}{\frac{1}{2}d_{00}^{L}+{\rm E}(a^{-1})}$
  $\bullet$ update ${\rm E}(a^{-1})\leftarrow\dfrac{1}{{\rm E}(\tau_{0}^{-1})+A^{-1}}$,

### v9

@@ Abstract
We have utilized the non-conjugate Variational Bayesian (VB) method for the problem of the sparse Poisson regression model. To provide approximate conjugacy in the model, the likelihood is approximated by a quadratic function, yielding conjugacy between the approximation component and the Gaussian prior on the regression coefficient. Three sparsity-enforcing priors (Laplace, Continuous Spike and Slab, and Bernoulli) are used for this problem. The proposed models are compared with each other, the associated MCMC models, and two frequentist sparse Poisson methods (LASSO and SCAD) to evaluate their estimation, prediction, and sparsity performance. In a simulation study, the proposed VB methods closely approximate the posterior parameter distribution while achieving significantly faster computation than benchmark MCMC methods. Using several benchmark count response data sets, the prediction performance of the proposed methods is evaluated in real-world applications.
@@ 1 Introduction
Although MCMC is the most widely used approach for approximating the posterior distribution in Bayesian inference, it incurs a high computational cost, particularly when the dimensionality of the regression parameters is large. There are a variety of methods for approximation of the posterior distribution with a lower computational cost, including integrated nested Laplace approximation (INLA- Rue et al., 2010), mean-field variational Bayes (VB- Blei et al., 2017), expectation propagation (EP- Minka, 2001, 2013), and message passing (MP- Winn et al., 2005). The mean-field VB is the most frequently used method in this category. The VB method is applied in many problems, especially in Bayesian regression models (see Šmídl and Quinn, 2006; Minka, 2000; Wand et al., 2011; Broderick et al., 2013; Drugowitsch, 2013; Goldsmith et al., 2011; Pham et al., 2013; Titsias and Lázaro-Gredilla, 2011; Sommerfelt et al., 2024; Berchuck et al., 2024; Anceschi et al., 2024, and the references therein). The VB method is also applied to the sparse linear regression models (Ormerod et al., 2017; Ray and Szabó, 2022; Titsias and Lázaro-Gredilla, 2011; Huang et al., 2016).
@@
The rest of the paper is organized as follows. In Section 2, the mean-field variational Bayes method is introduced. Three variational Bayesian sparse Poisson regression models are then proposed in Section 3. The derivation of the components, as well as the evidence lower bound (ELBO) for these methods, is given in Appendix A. In Section 4, a hard thresholding method is introduced to derive the final sparse estimates of the regression coefficients. Section 5 introduces the posterior predictive density function for predicting the response variable based on new covariate samples. Section 6 presents a simulation study for evaluating the estimation, prediction, and sparsity performance of the three proposed methods and comparing them with two frequentist sparse Poisson regression methods (LASSO and SCAD Poisson regression models) and the associated MCMC methods. A single replication of the simulation study is also used for a visual comparison of the MCMC empirical posteriors with their approximated counterparts. Some of the plots and additional analysis are given in supplementary material for the sake of brevity. Finally, in Section 7, the test set prediction performance of the proposed methods is compared with that of the frequentist alternatives, using 6 benchmark count response data sets. Some concluding remarks are given in Section 8.
@@ 2 Mean field variational Bayes
$q({\boldsymbol{\theta}})=\prod_{j=1}^{M}q({\boldsymbol{\theta}}_{j}),$
that is we assume ${\boldsymbol{\theta}}_{1},\ldots,{\boldsymbol{\theta}}_{M}$ to be independent, in the mean-field approximation.
@@ 3.1 Laplace prior
Figure 1: A directed acyclic graph (DAG) representing the hierarchical structure of Model (3). The data hierarchy is shown from top to bottom: observed data (squares) are generated from latent probabilistic processes, which are governed by model parameters (circles) and fixed hyperparameters (polygons). Arrows indicate conditional dependencies; for instance, the response variable $y$ depends on the parameters $\beta$ and $z$, illustrating the model’s core data-generating mechanism.
@@
Figure 3: Approximation error analysis for Taylor-expanded Poisson log-likelihood across different $\lambda$ ranges. The relative error in log-likelihood (as percentage) is shown against the percentage perturbation in $\log(\lambda)$. Solid lines represent mean errors, while dashed lines indicate the 2.5% and 97.5% quantiles from 1000 simulations.
@@
Since, the equality holds in (4) when $\xi=x$, we let $\xi_{i}={\rm E}_{q}(\mathbf{X}_{i}{\boldsymbol{\beta}})=\mathbf{X}_{i}{\color[rgb]{0,0,0}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(L)}}$, for $i=1,\ldots,n$, where $\color[rgb]{0,0,0}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(L)}$} is given in (6). Let $\boldsymbol{\xi}=(\xi_{1},\ldots,\xi_{n})^{\top}$, and define $M_{\boldsymbol{\xi}}=e^{\xi}(1-\xi)$ and $S_{X}^{\boldsymbol{\xi}}=\sum_{i=1}^{n}e^{\xi_{i}}\mathbf{X}_{i}\mathbf{X}_{i}^{\top}$.
To evaluate the precision of this approximation, a simulation study was conducted. For each of four different ranges of the Poisson rate parameter $\lambda$—specifically (1,10), (11,20), (21,30), and (31,40)—we generated 1000 datasets. Each dataset contained 100 observations, where the counts $y$ were drawn from a Poisson distribution with rates $\lambda$ drawn uniformly from the respective interval. The approximation error was assessed by introducing controlled perturbations to $\log(\lambda)$, varying the maximum perturbation size from 1% to 15%. For each scenario, the relative error between the true and the approximated log-likelihood was computed. The results, visualized in Figure 3, plot the percentage change in the log-likelihood against the percentage change in $\log(\lambda)$. Each panel displays the mean relative error (solid line) along with the 2.5% and 97.5% quantiles (dashed lines) across the simulations. The analysis reveals that the approximation error is generally small (often below 1%) for modest perturbations and smaller values of $\lambda$. However, both the mean error and its variability tend to increase with larger perturbation sizes and higher $\lambda$ ranges. This confirms that the Taylor approximation remains sufficiently accurate for practical use in VB inference for Poisson regression, provided the variational distribution does not induce excessive dispersion in $\log(\lambda)$, especially when the underlying rates are large.
@@
${\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(L)}=\left[S_{X}^{\boldsymbol{\xi}}+\text{diag}{\rm E}_{q}({\boldsymbol{\tau}}^{-1})\right]^{-1},\quad{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(L)}={\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(L)}\mathbf{X}^{\top}(\mathbf{y}-M_{\boldsymbol{\xi}}).$ (6)
Let $D^{(L)}_{{\boldsymbol{\beta}}}={\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(L)}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(L)}^{\top}+{\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(L)}$. Furthermore,
$q({\eta})=\text{Gamma}({\eta};{p+\nu-1},\delta+\frac{1}{2}\sum_{j=1}^{p-1}{\rm E}_{q}(\tau_{j})),$
and ${\rm E}_{q}({\eta})=\frac{{p+\nu-1}}{\delta+\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}_{q}(\tau_{j})}$, for $j=1,\ldots,p$,
$q(\tau_{j})=\text{Gen-Inv-Gauss}\Big(\tau_{j};\frac{1}{2},{\rm E}_{q}({\eta}),{\rm E}_{q}(\beta_{j}^{2})\Big),$
@@
${\rm E}_{q}(\tau_{j})=\frac{\sqrt{{\rm E}_{q}(\beta_{j}^{2})}K_{3/2}(\sqrt{{\rm E}_{q}({\eta}){\rm E}_{q}(\beta_{j}^{2})})}{\sqrt{{\rm E}_{q}({\eta})}K_{1/2}(\sqrt{{\rm E}_{q}({\eta}){\rm E}_{q}(\beta_{j}^{2})})},$
${\rm E}_{q}(\tau_{j}^{-1})=\frac{\sqrt{{\rm E}_{q}({\eta})}K_{3/2}(\sqrt{{\rm E}_{q}({\eta}){\rm E}_{q}(\beta_{j}^{2})})}{\sqrt{{\rm E}_{q}(\beta_{j}^{2})}K_{1/2}(\sqrt{{\rm E}_{q}({\eta}){\rm E}_{q}(\beta_{j}^{2})})}-\frac{1}{{\rm E}_{q}(\beta_{j}^{2})},$
@@
${\rm E}_{q}(\log\tau_{j})=\log\sqrt{\frac{{\rm E}_{q}(\beta_{j}^{2})}{{\rm E}_{q}({\eta})}}+\frac{\partial}{\partial t}\log K_{t}(\sqrt{{\rm E}_{q}({\eta}){\rm E}_{q}(\beta_{j}^{2})})|_{t=1/2}.$
@@
$q(\tau_{0})=\text{Inv-Gamma}(\tau_{0};1,\frac{1}{2}d_{00}^{(L)}+{\rm E}_{q}(a^{-1})),$
where ${d_{jj}^{(L)}}={\rm E}_{q}(\beta_{j}^{2})=(D^{(L)}_{{\boldsymbol{\beta}}})_{jj}$, and thus ${\rm E}_{q}(\tau_{0}^{-1})=\dfrac{1}{\frac{1}{2}d_{00}^{(L)}+{\rm E}_{q}(a^{-1})}$ and ${\rm E}_{q}(\log\tau_{0})=\log(\frac{1}{2}d_{00}^{(L)}+{\rm E}_{q}(a^{-1}))-\psi(1)$, where $\psi(\cdot)$ is the digamma function, and
$q(a)=\text{Inv-Gamma}(a;1,{\rm E}_{q}(\tau_{0}^{-1})+A^{-1}),$
and therefore ${\rm E}_{q}(a^{-1})=\dfrac{1}{{\rm E}_{q}(\tau_{0}^{-1})+A^{-1}}$ and ${\rm E}_{q}(\log a)=-\psi(1)+\log({\rm E}_{q}(\tau_{0}^{-1})+A^{-1}))$.
@@
[…]
${\color[rgb]{0,0,0}\quad+{\rm E}_{q}(\log p(\tau_{0}|a))+{\rm E}_{q}(\log p(a))-{\rm E}_{q}(\log q({\boldsymbol{\beta}}))\allowdisplaybreaks}$
${\color[rgb]{0,0,0}\quad-{\rm E}_{q}(\log q({\eta}))-{\rm E}_{q}(\log q(\tau_{0}))-{\rm E}_{q}(\log q(\tau))-{\rm E}_{q}(\log q(a))\allowdisplaybreaks}$
$=-M_{\boldsymbol{\xi}}^{T}(1+\mathbf{X}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(L)})-\frac{1}{2}\sum_{i=1}^{n}\xi_{i}^{2}e^{\xi_{i}}-\frac{1}{2}tr(S_{X}^{\boldsymbol{\xi}}D_{{\boldsymbol{\beta}}}^{(L)})+\mathbf{y}^{\top}\mathbf{X}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(L)}\allowdisplaybreaks$
$\quad-\frac{1}{2}{\rm E}_{q}(\log\tau_{0})-\frac{1}{2}{\rm E}_{q}(\tau_{0}^{-1}){\rm E}_{q}(\beta_{0}^{2})-\dfrac{\delta({p+\nu-1})}{\delta+\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}_{q}(\tau_{j})}\allowdisplaybreaks$
$\quad-2\log(A^{-1}+{\rm E}_{q}(\tau_{0}^{-1}))-\frac{1}{2}\log(d_{00}^{(L)}/2+{\rm E}_{q}(a^{-1}))\allowdisplaybreaks$
$\quad-{\rm E}_{q}(a^{-1}){\rm E}_{q}(\tau_{0}^{-1})-\dfrac{A^{-1}}{A^{-1}+E(\tau_{0}^{-1})}+\frac{1}{2}\log|{\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(L)}|\allowdisplaybreaks$
$\quad-\frac{1}{4}{\sum_{j=1}^{p-1}}\frac{{\rm E}_{q}({\eta})}{{d_{jj}^{(L)}}}+{\sum_{j=1}^{p-1}}\log K_{1/2}(\sqrt{{\rm E}_{q}({\eta}){d_{jj}^{(L)}}})+\text{Const.}$ (7)
@@
  $\bullet$ update ${\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(L)}\leftarrow\left[S_{X}^{\boldsymbol{\xi}}+\text{diag}{\rm E}_{q}({\boldsymbol{\tau}}^{-1})\right]^{-1},$ and ${\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(L)}\leftarrow{\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(L)}\mathbf{X}^{\top}(\mathbf{y}-M_{\boldsymbol{\xi}}),$
@@
  $\bullet$ update ${\rm E}_{q}(\beta_{j}^{2})={d_{jj}^{(L)}}\leftarrow{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(L)j}^{2}+{\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(L)jj}$, for $j=0,\ldots,p-1$,
  $\bullet$ update ${\rm E}_{q}({\eta})\leftarrow\frac{{p+\nu-1}}{\delta+\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}_{q}(\tau_{j})}$
  $\bullet$ update ${\rm E}_{q}(\tau_{j})\leftarrow\frac{\sqrt{{\rm E}_{q}(\beta_{j}^{2})}K_{3/2}(\sqrt{{\rm E}_{q}({\eta}){\rm E}_{q}(\beta_{j}^{2})})}{\sqrt{{\rm E}_{q}({\eta})}K_{1/2}(\sqrt{{\rm E}_{q}({\eta}){\rm E}_{q}(\beta_{j}^{2})})},$ for $j=1,\ldots,p-1$,
  $\bullet$ update ${\rm E}_{q}(\tau_{j}^{-1})\leftarrow\frac{\sqrt{{\rm E}_{q}({\eta})}K_{3/2}(\sqrt{{\rm E}_{q}({\eta}){\rm E}_{q}(\beta_{j}^{2})})}{\sqrt{{\rm E}_{q}(\beta_{j}^{2})}K_{1/2}(\sqrt{{\rm E}_{q}({\eta}){\rm E}_{q}(\beta_{j}^{2})})}-\frac{1}{{\rm E}_{q}(\beta_{j}^{2})},$ for $j=1,\ldots,p-1$,
  $\bullet$ update ${\rm E}_{q}(\tau_{0}^{-1})\leftarrow\dfrac{1}{\frac{1}{2}d_{00}^{(L)}+{\rm E}_{q}(a^{-1})}$
  $\bullet$ update ${\rm E}_{q}(a^{-1})\leftarrow\dfrac{1}{{\rm E}_{q}(\tau_{0}^{-1})+A^{-1}}$,